Y = In ( e^x + e^-x ) ; a = In^3
1
Expert's answer
2021-09-01T11:35:01-0400

The question is incomplete but we can find derivative here.

"y=ln(e^x+e^{-x})
\\\Rightarrow \dfrac{dy}{dx}=\dfrac{d}{dx}(ln(e^x+e^{-x}))
\\\Rightarrow \dfrac{dy}{dx}=\dfrac{1}{e^x+e^{-x}}\times \dfrac{d}{dx}(e^x+e^{-x})
\\\Rightarrow \dfrac{dy}{dx}=\dfrac{1}{e^x+e^{-x}}\times(e^x-e^{-x})
\\\Rightarrow \dfrac{dy}{dx}=\dfrac{e^x-e^{-x}}{e^x+e^{-x}}"